# Eternal Universes

### v1

@@ V.1 Loitering eternal universe
It is possible to construct a loitering eternal universe by smoothly gluing together a long Minkowski past phase to an inflating phase via a smooth bump function. This blueprint was first utilized in our previous work Lesnefsky et al. (2023b) to build a loitering model which is complete.
For this example we take the scale factor to be constant for $t<0$, so that $H=0\,\forall t<0$, and the spacetime is Minkowski. For $t>1$, the scale factor is exponential, $H$ is constant and the spacetime is qausi-de Sitter. A smooth bump function is used to piece the two regimes together in the regime $t\in[0,1]$.
@@
Clearly the model is eternal and complete. As with previous examples, the NEC is violated when $\dot{H}>0$, and inflationary by Defn. 1 with $\ddot{a}>0$ for $t>0$. The ANEC Eq. 1.1 is satisfied. The Minkowski period with $H=0$ satisfies the bouncing conditions with $\dot{a}=0$ and with non-trivial scale factor overall, of Defn. 2. The model has an accelerating de Sitter phase at late times with $H=const$.
@@ V.2 Quasi-loitering eternal universe
For our final example of an eternal universe we construct what we refer to as a quasi-loitering universe in that the scale factor is always expanding, although initially at an exponentially small (yet accelerated) rate before entering an increased accelerated inflationary phase. The situation at early time is reminiscent of the so-called Galileon Genesis model Creminelli et al. (2010), in which a universe that is asymptotically Minkowski in the past expands with $\dot{H}>H^{2}$, violating the NEC with an increasing energy density $\rho$; although, in the current model this phase naturally comes to an end. Eventually the inflationary phase transitions, the energy density decreases and the universe enters another period of extremely slow (negative accelerated) non-inflationary expansion. The quasi-loitering scale factor is:
@@
In order to achieve geodesic completeness we require sufficient lifting of the tanh function, requiring $c>1$. The Hubble parameter is:
@@
[…]
@@
Figure 11: Evolutionary behavior for Eq. 5.34. The scale factor (red). The Hubble parameter (green). The co-moving Hubble radius suppressed by 1000 (blue). Model parameters: $a_{0}=\alpha=1$$c=2$.
@@
Figure 13: Energy conditions from Eq. V.2. Plot of energy density $\rho$ (red), $\rho+p$ (yellow), $|p|$ (blue) and $\rho+3p$ (green). Model parameters: $a_{0}=\alpha=1$, $c=2$.
At early times, all of the classical energy conditions are violated. At late times the WEC, NEC and SEC are all satisfied. Only the DEC is mildly violated at late times; however, there exists a brief period for small positive $t$, during which all of the energy conditions are satisfied. This period lasts from $t=0$ when the SEC (green curve) starts to be satisfied until the moment the blue curve grows larger than the red curve where the DEC begins to be violated. This healthy period can be past extended to accommodate an inflationary phase where only the SEC is violated (starting where NEC is valid and yellow curve intersects the horizontal line). The inflationary period naturally ends when the SEC starts to be obeyed (green curve intersects the horizontal line). For the ANEC, we evaluate Eq. 1.1
[…]
for our model parameters. Hence, the ANEC is satisfied.
@@ VI Discussion
In General Relativity, every smooth, non-constant scale factor a(t) of a geodesically complete FRW spacetime must violate the NEC for at least some period of time.
The above is obviously true for all bouncing FRW cosmological models. As is well-known, in a flat $k=0$ (or hyperbolic $k=-1$) FRW spacetime, the NEC must be violated in order to achieve a cosmological bounce Hawking and Penrose (1970). During a bounce through a local minimum, the universe transitions from a contracting phase, $H<0$, to an expanding phase, $H>0$. This transition inherently requires that at the point where the contraction halts and expansion begins, the derivative of the Hubble parameter, $\dot{H}$, must be positive. Since $\dot{H}=-4\pi G(\rho+p)$, we must have $\rho+p<0$, signalling violation of the NEC. Thus, any such spacetime which exhibits a bounce, or bounces, for any part of it’s history must violate NEC. Likewise any super-inflationary models (by definition) have $\dot{H}>0$ and violate NEC. With non-zero curvature $k$, one can produce a bounce without violating NEC since, $\dot{H}=-4\pi G(\rho+p)+k/a^{2}$, can become positive at the bounce due to positive $k$; however, such universes are closed if the matter sector obeys NEC. While the NEC holds, $\ddot{a}/a=H^{2}+\dot{H}=-4\pi G(\rho/3+p)$, and $H=0$ at the bounce, $\ddot{a}/a>0$ and the SEC is violated. Curvature bounces are discussed in detail in Starobinskii (1978); Graham et al. (2014).
@@
and is compatible with Eq. 4.23. Applied to the scale factor Eq. 3.11, the odd powered polynomial terms are negative as $t\rightarrow-\infty$, and effectively cancel the positive even powered terms ultimately creating a “bounce point at infinity” in accordance with Defns. 2,3 at $t\rightarrow-\infty$, with constant scale factor $a=c$ and with $H=0$. Thus, our eternal inflating model Eq. 3.11 is also a bouncing cosmology par Defn.2. To elucidate our discussion we remind the reader of the general polynomial bounce discussed in §IV.2 and Eqs. 4.23 and 4.24. In essence the bounce point of a finite series polynomial, such as Eq. 4.24, is in this case translated to $t\rightarrow-\infty$, where
@@
In General Relativity, every smooth, non-constant scale factor a(t) of a geodesically complete FRW spacetime must experience at least one bounce.
Combining this with Cors.4,6 and Conj.5, we formulate a provocative composite conjecture:
@@
Every geodesically complete eternal spacetime which admits a neighborhood isometric to a GFRW with a non-constant scale factor will inflate for some time and bounce at least once.
@@
In our discussion of a fully geodesically complete spacetime that exists for an eternity, we encounter a tautological paradigm shift: the key focus moves from categorizing spacetimes as “inflationary” or “bouncing” to emphasizing their geodesic completeness. We discover from Conj. 9, although both inflation and cosmological bounces play crucial roles in forming a geodesically complete spacetime, they are often transient stages in the vast timeline of an eternal cosmos.  [^14] An observer inhabiting a decelerating universe, who holds completeness as a foundational principle, would predict the existence of a prior inflationary period by Cor. 4, and even anticipate the likelihood of a future phase of acceleration.
@@ Appendix B Theorem 1 Item 3 Proof
Considering this, a periodic spacelike trajectory will have $d\pi_{\mathbb{R}^{1}_{1}}\dot{\gamma}=\frac{dt}{d\lambda}\frac{\partial}{\partial t}$, partially future pointing and partially past pointing so that the two distinct past (future) directed integrals coincide. For the case of unbounded trajectories, if the null integral diverges so will the two past (future) pointing spacelike integrals, for the reason mentioned above. Thus, for spacelike completeness, one must be both past (future) null complete - along with the achronal case which holds by assumption - so full null completeness must hold. The null case of the integral of item (2) bounds the non-infinite $a$ spacelike case from below. Thus, if the GFRW is past and / or future null complete then - in the case of a non-infinite $a$ the GFRW is spacelike complete and the first portion of item (3) is proven. For completeness, we include the case where $a$ potentially diverges and becomes infinite for a finite $t\in\mathbb{R}$. In this case a “big rip” boundary will form and a geodesic will encounter it within finite parameter.
@@ C.1 Definition
$\frac{d^{2}x^{\sigma}}{d\lambda^{2}}+\Gamma^{\sigma}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}=0\,.$ (3.46)
@@ C.2 Geodesics in GFRWs
$\nabla_{\frac{dx^{\nu}}{d\lambda}\frac{\partial}{\partial x^{\nu}}}\frac{dx^{\mu}}{d\lambda}\frac{\partial}{\partial x^{\mu}}=\left(\frac{d^{2}x^{\mu}}{d\lambda^{2}}+\Gamma^{\mu}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}\right)\frac{\partial}{\partial x^{\mu}}=0$ (3.47)
@@
$\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}=-g_{\Sigma}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)\cdot\left(a\left(t\left(\lambda\right)\right)\right)\cdot\frac{da}{dt}$ (3.48)
$\nabla_{\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)}\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)=-\frac{2}{a\left(t\left(\lambda\right)\right)}\frac{da}{dt}\frac{dt}{d\lambda}\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}$ (3.49)
respectively. In particular, Eq. 3.49 in $\Sigma$ is not quite totally geodesic, but is a pregeodesic because the acceleration is proportional to $\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}$ and thus a geodesic reparameterization exists “boosting” away the acceleration the affine parameter $\lambda$. The image of Eq. 3.49 is the image of a geodesic. We now invoke the completeness assumption of $\left(\Sigma,g_{\Sigma}\right)$, yielding $x^{k}\left(\lambda\right)$ is defined $\forall\,\mathbb{R}$. Finally, if $x^{\mu}\left(\lambda\right)$ fails to be complete, it must be due to Eq. 3.48.
@@ C.3 Derivation of Warped Product Geodesics
In fact, Eqs. 3.48 and 3.49 can be generalized to a warped product of any semi-Riemannian manifolds. The familiar geodesic equation Eq. 3.46 can be distilled into the generalized form of Eqs. 3.48 and 3.49. In the following we use $f$ for the scale factor. As in O’Neill (1983) Chapter 7 in a warped product of any two semi-Riemannian manifolds $B\times_{f}F$ the first manifold is the leaf manifold $\left(B,\tilde{g}_{B}\right)$ corresponding to $\mathbb{R}^{1}_{1}$ and the second manifold is the fiber $\left(F,\tilde{g}_{F}\right)$ corresponding to $\left(\Sigma,g_{\Sigma}\right)$. Additionally, consider a chart $\left(U_{B},\{t^{k}\}\right)$ of $B$ and a chart $\left(U_{F},\{x^{\ell}\}\right)$. The warped product metric is a block diagonal form $g_{B}+f^{2}\left(t^{k}\right)g_{F}$. All factors of $f\left(t^{k}\right)$ will be explicitly written and terms with a “tilde” are viewed to be elements of the constituent spaces and naked (no tilde) terms are elements of the warped product $B\times_{f}F$, for example the pullback $g_{B}=d\pi_{B}^{*}\tilde{g}_{B}$. Finally, the geodesic equation is calculated over curve $t^{k}\left(\lambda\right)\otimes x^{\ell}\left(\lambda\right)$.
@@
$\Gamma^{t^{k}}_{t^{\ell}t^{m}}$ $=\tilde{\Gamma}^{t^{k}}_{t^{\ell}t^{m}}$ $\Gamma^{t^{k}}_{x^{\ell}x^{m}}$ $=-\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}f\left(t^{k}\right)\mathrm{grad}_{\tilde{g}_{B}}f$ $\Gamma^{t^{k}}_{t^{\ell}x^{m}}$ $=0$
@@
00 0 $=\frac{d^{2}x^{\sigma}}{d\lambda^{2}}+\Gamma^{\sigma}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}$
@@
$=\frac{d^{2}t^{k}}{d\lambda^{2}}+\tilde{\Gamma}^{t^{k}}_{t^{a}t^{m}}\frac{dt^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}-\left(\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}\right)\cdot\left(f\left(t^{k}\right)\mathrm{grad}_{\tilde{g}_{B}}f\right)$
$\qquad+\frac{d^{2}x^{\ell}}{d\lambda^{2}}+\tilde{\Gamma}^{x^{\ell}}_{x^{a}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}+\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\left(\tilde{g}_{F}\right)^{x^{\ell}x^{a}}\left(\tilde{g}_{F}\right)_{x^{a}x^{k}}\frac{dx^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}$ (3.50)
@@
$\frac{d^{2}t^{k}}{d\lambda^{2}}+\tilde{\Gamma}^{t^{k}}_{t^{a}t^{m}}\frac{dt^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}=\left(\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}\right)\cdot\left(f\left(t^{k}\right)\mathrm{grad}_{\tilde{g}_{B}}f\right)$ (3.51)
$\frac{d^{2}x^{\ell}}{d\lambda^{2}}+\tilde{\Gamma}^{x^{\ell}}_{x^{a}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}=-\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\delta^{x^{\ell}}_{x^{k}}\frac{dx^{k}}{d\lambda}\frac{dt^{m}}{d\lambda}$ (3.52)
are recovered. In order to eventually compare Eqs. 3.51, 3.52 to Eqs. 3.48, 3.49 in the case of a GFRW one can consolidate using formal notation  [^16] A pedantic reader may have noticed the “free” index of $t^{\ell}$ on the right hand side of Eq. 3.53. This is an unfortunate thorn of mixing “component” and “geometer” notation. Strictly – and in most geometric texts such as O’Neill (1983); Beem et al. (1996); Peterson (2006) the projection of the geodesic onto $B$ is written $\tau:J\rightarrow B$ and the function written as the composition $\left(f\circ\tau\right)\left(\lambda\right)$. While being precise, this minimizes the “component” moiety of the expression which we are trying to draw attention to. A similar occurrence appears in Eq. 3.54. We have been a bit caviler with the notation in order to show the relationship between the two notations - we hope the reader will forgive this indulgence of an unbalanced “free” index.
$\nabla_{\left(\frac{dt^{\ell}}{d\lambda}\frac{\partial}{\partial t^{\ell}}\right)}\left(\frac{dt^{\ell}}{d\lambda}\frac{\partial}{\partial t^{\ell}}\right)=\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{m}}{d\lambda}\frac{\partial}{\partial x^{m}}\right)\cdot\left(f\left(t^{\ell}\right)\mathrm{grad}_{\tilde{g}_{B}}f\right)$ (3.53)
$\nabla_{\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)}\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)=-\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\frac{dt^{m}}{d\lambda}\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}$ (3.54)
In the case of a GFRW, $\left(B,\tilde{g}_{B}\right)=\mathbb{R}^{1}_{1}$ and $\left(F,\tilde{g}_{F}\right)=\left(\Sigma,g_{\Sigma}\right)$; with respect to Eqs. 3.53, 3.54 the aforementioned Eqs. 3.48, 3.49 are recovered, as desired.

### v2

@@ V.1 Loitering eternal universe
It is possible to construct a loitering eternal universe by smoothly gluing together a long Minkowski past phase to an inflating phase via a smooth bump function. This blueprint was first utilized in our previous work Lesnefsky et al. (2023) to build a loitering model which is complete.
For this example we take the scale factor to be constant for $t<0$, so that $H=0\,\forall t<0$, and the spacetime is Minkowski. For $t>1$, the scale factor is exponential, $H$ is constant and the spacetime is quasi-de Sitter. A smooth bump function is used to piece the two regimes together in the regime $t\in[0,1]$.
@@
Clearly the model is eternal and complete. As with previous examples, the NEC is violated when $\dot{H}>0$, and inflationary by Defn. 1 with $\ddot{a}>0$ for $t>0$. The Minkowski period with $H=0$ satisfies the bouncing conditions with $\dot{a}=0$ and with non-trivial scale factor overall, of Defn. 2. The model has an accelerating de Sitter phase at late times with $H=const$. The “phantom-like” physics required to push the universe from the static $H=0$ plateau into the de Sitter phase imprints an overall negative ANEC integral.
@@ V.2 Quasi-loitering eternal universe
For our final example of an eternal universe we construct what we refer to as a quasi-loitering universe–in that the scale factor is always expanding, although initially at an exponentially small (yet accelerated) rate before entering an increased accelerated inflationary phase. The situation at early time is reminiscent of the so-called Galileon Genesis model Creminelli et al. (2010), in which a universe that is asymptotically Minkowski in the past expands with $\dot{H}>H^{2}$, violating the NEC with an increasing energy density $\rho$; although, in the current model this phase naturally comes to an end. Eventually the inflationary phase transitions, the energy density decreases and the universe enters another period of extremely slow (negative accelerated) non-inflationary expansion. The quasi-loitering scale factor is:
@@
In order to achieve geodesic completeness we require sufficient lifting of the tanh function, requiring $c>a_{0}$. The Hubble parameter is:
@@
$\int^{t}a\left(\zeta\right)d\zeta=ct+a_{0}\alpha\log\left(\cosh\left(\frac{t}{\alpha}\right)\right)\,.$ (5.48)
@@
Figure 11: Evolutionary behavior for Eq. 5.44. The scale factor (red). The Hubble parameter (green) magnified by an order of magnitude. The co-moving Hubble radius suppressed by an order of magnitude (blue). Model parameters: $a_{0}=\alpha=1$$c=2$.
@@
Figure 13: Energy conditions from Eq. V.2. Plot of energy density $\rho$ (blue), $\rho+p$ (yellow), $|p|$ (red) and $\rho+3p$ (green). Legend quantities are rendered dimensionless by the factor $\alpha^{2}/M^{2}_{pl}$. Model parameters: $a_{0}=\alpha=1$, $c=2$.
At early times, all of the classical energy conditions are violated. At late times the WEC, NEC and SEC are all satisfied. Only the DEC is mildly violated at late times; however, there exists a brief period for small positive $t$, during which all of the energy conditions are satisfied. This period lasts from $t=0$ when the SEC (green curve) starts to be satisfied until the moment the red curve grows larger than the blue curve where the DEC begins to be violated. This healthy period can be past extended to accommodate an inflationary phase where only the SEC is violated (starting where NEC is valid and yellow curve intersects the horizontal line). The inflationary period naturally ends when the SEC starts to be obeyed (green curve intersects the horizontal line). For the ANEC, we evaluate Eq. 1.1
$\int_{-\infty}^{\infty}\left(\rho+p\right)d\lambda$ $=\frac{2}{\alpha a_{0}}\left(\frac{c}{a_{0}\tanh\left(\frac{t}{\alpha}\right)+c}+\log\left(a_{0}\sinh\left(\frac{t}{\alpha}\right)+c\cosh\left(\frac{t}{\alpha}\right)\right)\right$.
$\qquad\qquad\left.\left.-\log\left(\cosh\left(\frac{t}{\alpha}\right)\right)-\frac{a_{0}^{2}}{2\left(a_{0}\sinh\left(\frac{t}{\alpha}\right)+c\cosh\left(\frac{t}{\alpha}\right)\right)^{2}}-1\right)\right|_{-\infty}^{\infty}$
$\leq 0\,.$ (5.50)
Hence, the ANEC is violated.
@@ VI Discussion
In General Relativity, every smooth, non-constant scale factor a(t) of a geodesically complete, flat FRW spacetime must violate the NEC during at least some period of time.
For example, it is well known in a flat $k=0$ (or hyperbolic k=1k=-1) FRW spacetime, the NEC must be violated in order to achieve a cosmological bounce Hawking and Penrose (1970). During a bounce through a local minimum, the universe transitions from a contracting phase, $H<0$, to an expanding phase, $H>0$. This transition inherently requires that at the point where the contraction halts and expansion begins, the derivative of the Hubble parameter, $\dot{H}$, must be positive. Since $\dot{H}=-4\pi G(\rho+p)$, we must have $\rho+p<0$, signalling violation of the NEC. Thus, any such spacetime which exhibits a bounce, or bounces, for any part of it’s history must violate NEC. Likewise any super-inflationary models (by definition) have $\dot{H}>0$ and violate NEC. With non-zero curvature $k$, one can produce a bounce without violating NEC since, $\dot{H}=-4\pi G(\rho+p)+k/a^{2}$, can become positive at the bounce in the case of positive $k$; however, such universes are closed if the matter sector obeys NEC and SEC. While the NEC holds, $\ddot{a}/a=H^{2}+\dot{H}=-4\pi G(\rho/3+p)$, and $H=0$ at the bounce, $\ddot{a}/a>0$ and the SEC is violated. Curvature bounces are discussed in detail in Starobinskii (1978); Graham et al. (2014); Burwig and Easson (2025).
@@
and is compatible with Eq. 4.23. Applied to the scale factor Eq. 3.11, the odd powered polynomial terms are negative as $t\rightarrow$-$\infty$, and effectively cancel the positive even powered terms ultimately creating a “bounce point at infinity” in accordance with Defns. 2,3 at $t\rightarrow$-$\infty$, with constant scale factor $a=c$ and with $H=0$. Thus, our eternal inflating model Eq. 3.11 is also a bouncing cosmology par Defn.2. To elucidate our discussion we remind the reader of the general polynomial bounce discussed in §IV.2 and Eqs. 4.23 and 4.34. In essence the bounce point of a finite series polynomial, such as Eq. 4.34, is in this case translated to $t\rightarrow$-$\infty$, where
@@
In General Relativity, every smooth, non-constant scale factor a(t) of a geodesically complete FRW spacetime must experience at least one bounce, where said bounce may be at infinity.
Combining this with Cor.4 we formulate a provocative composite conjecture:
@@
Every geodesically complete eternal spacetime which admits a neighborhood isometric to a GFRW with a non-constant scale factor will inflate for some time and bounce at least once, where said bounce may be at infinity.
@@
In our discussion of a fully geodesically complete spacetime that exists for an eternity, we encounter a tautological paradigm shift: the key focus moves from categorizing spacetimes as “inflationary” or “bouncing” to emphasizing their geodesic completeness. We discover from Conj. 8, although both inflation and cosmological bounces play crucial roles in forming a geodesically complete spacetime, they are often transient stages in the vast timeline of an eternal cosmos.  [^15] An observer inhabiting a decelerating universe, who holds completeness as a foundational principle, would predict the existence of a prior inflationary period by Cor. 4, and even anticipate the likelihood of a future phase of acceleration.
@@ Appendix B Theorem 1 Item 3 Proof
Considering this, a periodic spacelike trajectory will have $d\pi_{\mathbb{R}^{1}_{1}}\dot{\gamma}=\frac{dt}{d\lambda}\frac{\partial}{\partial t}$, partially future pointing and partially past pointing so that the two distinct past (future) directed integrals coincide. For the case of unbounded trajectories, if the null integral diverges so will the two past (future) pointing spacelike integrals, for the reason mentioned above. Thus, for spacelike completeness, one must be both past (future) null complete - along with the achronal case which holds by assumption - so full null completeness must hold. The null case of the integral of item (2) bounds the non-infinite $a$ spacelike case from below. Thus, if the GFRW is past and / or future null complete then - in the case of a non-infinite $a$ - the GFRW is spacelike complete and the first portion of item (3) is proven. For completeness, we include the case where $a$ potentially diverges and becomes infinite for a finite $t\in\mathbb{R}$. In this case a “big rip” boundary will form and a geodesic will encounter it within finite parameter.
@@ C.1 Definition
$\frac{d^{2}x^{\sigma}}{d\lambda^{2}}+\Gamma^{\sigma}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}=0\,.$ (C.56)
@@ C.2 Geodesics in GFRWs
$\nabla_{\frac{dx^{\nu}}{d\lambda}\frac{\partial}{\partial x^{\nu}}}\frac{dx^{\mu}}{d\lambda}\frac{\partial}{\partial x^{\mu}}=\left(\frac{d^{2}x^{\mu}}{d\lambda^{2}}+\Gamma^{\mu}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}\right)\frac{\partial}{\partial x^{\mu}}=0$ (C.57)
@@
$\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}=-g_{\Sigma}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)\cdot\left(a\left(t\left(\lambda\right)\right)\right)\cdot\frac{da}{dt}$ (C.58)
$\nabla_{\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)}\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)=-\frac{2}{a\left(t\left(\lambda\right)\right)}\frac{da}{dt}\frac{dt}{d\lambda}\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}$ (C.59)
respectively. In particular, Eq. C.59 in $\Sigma$ is not quite totally geodesic, but is a pregeodesic because the acceleration is proportional to $\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}$ and thus a geodesic reparameterization exists “boosting” away the acceleration the affine parameter $\lambda$. The image of Eq. C.59 is the image of a geodesic. We now invoke the completeness assumption of $\left(\Sigma,g_{\Sigma}\right)$, yielding $x^{k}\left(\lambda\right)$ is defined $\forall\,\mathbb{R}$. Finally, if $x^{\mu}\left(\lambda\right)$ fails to be complete, it must be due to Eq. C.58.
@@ C.3 Derivation of Warped Product Geodesics
In fact, Eqs. C.58 and C.59 can be generalized to a warped product of any semi-Riemannian manifolds. The familiar geodesic equation Eq. C.56 can be distilled into the generalized form of Eqs. C.58 and C.59. In the following we use $f$ for the scale factor. As in O’Neill (1983) Chapter 7 in a warped product of any two semi-Riemannian manifolds $B\times_{f}F$ the first manifold is the leaf manifold $\left(B,\tilde{g}_{B}\right)$ - corresponding to $\mathbb{R}^{1}_{1}$ - and the second manifold is the fiber $\left(F,\tilde{g}_{F}\right)$ - corresponding to $\left(\Sigma,g_{\Sigma}\right)$. Additionally, consider a chart $\left(U_{B},\{t^{k}\}\right)$ of $B$ and a chart $\left(U_{F},\{x^{\ell}\}\right)$. The warped product metric is a block diagonal form $g_{B}+f^{2}\left(t^{k}\right)g_{F}$. All factors of $f\left(t^{k}\right)$ will be explicitly written and terms with a “tilde” are viewed to be elements of the constituent spaces and naked (no tilde) terms are elements of the warped product $B\times_{f}F$, for example the pullback $g_{B}=d\pi_{B}^{*}\tilde{g}_{B}$. Finally, the geodesic equation is calculated over curve $t^{k}\left(\lambda\right)\oplus x^{\ell}\left(\lambda\right)$.
@@
$\Gamma^{t^{k}}_{t^{\ell}t^{m}}$ $=\tilde{\Gamma}^{t^{k}}_{t^{\ell}t^{m}}$ $\Gamma^{t^{k}}_{x^{\ell}x^{m}}$ $=-\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}f\left(t^{k}\right)\left(\mathrm{grad}_{\tilde{g}_{B}}f\right)^{t^{k}}$ $\Gamma^{t^{k}}_{t^{\ell}x^{m}}$ $=0$
@@
where $\left(\mathrm{grad}_{\tilde{g}_{B}}f\right)^{t^{k}}$ is the $t^{k}$ component of the gradient, and the scale factor of $f:B\rightarrow\mathbb{R}$ is written $f\left(t^{k}\right)$.
@@
$0$ $=\frac{d^{2}x^{\sigma}}{d\lambda^{2}}+\Gamma^{\sigma}_{\alpha\beta}\frac{dx^{\alpha}}{d\lambda}\frac{dx^{\beta}}{d\lambda}$
@@
$=\frac{d^{2}t^{k}}{d\lambda^{2}}+\tilde{\Gamma}^{t^{k}}_{t^{a}t^{m}}\frac{dt^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}-\left(\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}\right)\cdot\left(f\left(t^{k}\right)\left(\mathrm{grad}_{\tilde{g}_{B}}f\right)^{t^{k}}\right)$
$\qquad+\frac{d^{2}x^{\ell}}{d\lambda^{2}}+\tilde{\Gamma}^{x^{\ell}}_{x^{a}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}+\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\left(\tilde{g}_{F}\right)^{x^{\ell}x^{a}}\left(\tilde{g}_{F}\right)_{x^{a}x^{k}}\frac{dx^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}$ (C.60)
@@
$\frac{d^{2}t^{k}}{d\lambda^{2}}+\tilde{\Gamma}^{t^{k}}_{t^{a}t^{m}}\frac{dt^{a}}{d\lambda}\frac{dt^{m}}{d\lambda}=\left(\left(\tilde{g}_{F}\right)_{x^{\ell}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}\right)\cdot\left(f\left(t^{k}\right)\left(\mathrm{grad}_{\tilde{g}_{B}}f\right)^{t^{k}}\right)$ (C.61)
$\frac{d^{2}x^{\ell}}{d\lambda^{2}}+\tilde{\Gamma}^{x^{\ell}}_{x^{a}x^{m}}\frac{dx^{a}}{d\lambda}\frac{dx^{m}}{d\lambda}=-\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\delta^{x^{\ell}}_{x^{k}}\frac{dx^{k}}{d\lambda}\frac{dt^{m}}{d\lambda}$ (C.62)
are recovered. In order to eventually compare Eqs. C.61, C.62 to Eqs. C.58, C.59 in the case of a GFRW one can consolidate using formal notation  [^17] A pedantic reader may have noticed the “free” index of $t^{\ell}$ on the right hand side of Eq. C.63. This is an unfortunate thorn of mixing “component” and “geometer” notation. Strictly – and in most geometric texts such as O’Neill (1983); Beem et al. (1996); Peterson (2006) the projection of the geodesic onto $B$ is written $\tau:J\rightarrow B$ and the function written as the composition $\left(f\circ\tau\right)\left(\lambda\right)$. While being precise, this minimizes the “component” moiety of the expression which we are trying to draw attention to. A similar occurrence appears in Eq. C.64. We have been a bit caviler with the notation in order to show the relationship between the two notations - we hope the reader will forgive this indulgence of an unbalanced “free” index.
$\nabla_{\left(\frac{dt^{\ell}}{d\lambda}\frac{\partial}{\partial t^{\ell}}\right)}\left(\frac{dt^{\ell}}{d\lambda}\frac{\partial}{\partial t^{\ell}}\right)=\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{m}}{d\lambda}\frac{\partial}{\partial x^{m}}\right)\cdot\left(f\left(t^{\ell}\right)\mathrm{grad}_{\tilde{g}_{B}}f\right)$ (C.63)
$\nabla_{\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)}\left(\frac{dx^{\ell}}{d\lambda}\frac{\partial}{\partial x^{\ell}}\right)=-\frac{2}{f\left(t^{m}\right)}\frac{\partial f}{\partial t^{m}}\frac{dt^{m}}{d\lambda}\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}$ (C.64)
In the case of a GFRW, $\left(B,\tilde{g}_{B}\right)=\mathbb{R}^{1}_{1}$ and $\left(F,\tilde{g}_{F}\right)=\left(\Sigma,g_{\Sigma}\right)$; with respect to Eqs. C.63, C.64 the aforementioned Eqs. C.58, C.59 are recovered, as desired.